# Bi-parental Reproduction may Enhance Species Sustainability by Conserving Shared Parental Traits more Faithfully than Mono-parental Reproduction

### v3

@@ 3.2 Reproduction methods
Below we describe one reproduction step, termed a transition in each of the methods.
Mono-parental Reproduction
@@
    Denote the current population as being the set $c$ and create an empty set $c^{\prime}$.
@@
      $p_{i}=\frac{1}{k}\sum_{\ell=1}^{k}x_{i_{\ell}}~{}~{};$
      this choice of reproduction step simplifies the model while keeping it intuitive and not limiting the claims; it can be generalized in future analyses.
@@
    After completing $n$ iterations: Mark the current population as being the set $c^{\prime}$ and discard $c$.
Bi-parental Reproduction
@@
    Denote the current population as being the set $c$ and create an empty set $c^{\prime}$.
@@
      Pick a random pair of two separate individuals, $e_{u},e_{v}\in c,~{}u\neq v,1\leq u,v\leq n$ out of the $\frac{n(n-1)}{2}$ possible pairs in $c$. The pair $e_{u},e_{v}$ will be the joint parents of the child created in this iteration; in separate iterations these two individuals may participate in other pairs or again in the very same pair.
@@
    After completing $n$ iterations: Mark the current population as being the set $c^{\prime}$ and discard $c$.
As stated earlier, neither MR nor BR employ considerations of Natural Selection and classical fitness metrics like an individual’s probability of reproduction and its expected number of offspring over its life time, and BR does not make use of mating types or of mate selection.
@@ 3.3 Model realization example
Below we illustrate the above formal model with an example realization in an imaginary species of intelligent and capable animals. One of the traits that sustains this species is their ability to build every year a mud hill of “exactly” a certain height, as traps for some creature that they feed upon. The way they remember the right height is that each mature individual carries with them a bunch of sticks that they got from their parents. The individual builds each mud hill so that its height is close to the mean length of its set of sticks. Old mud hills from the previous year are not available for comparison, as they have been washed away by the rains. When young individuals mature, they leave their parents to look for new territory. Before they leave, the parents (one or two of them, as the case may be) equip the young with a new bunch of sticks, cut as best as they can to the mean length of the parents’ bunch. In the case of two parents’ each one gives the child half the number of sticks based on that parent’s set. The parents cannot give their own sticks to the offspring, as they have to keep their own sticks for several more years. The individuals are solitary most of the time, so each one needs their own set of sticks. Individuals of the species use sets of sticks for this purpose, rather than just one, as the sticks may break or get lost. In addition, the parent’s skill and the available material limit the precision of each cut stick; having several sticks, some too short and some too long, is a way for conveying the height of the mud hill that is desired by the individual preparing the new set.
@@ 3.4 A numerical example
    Parameter Setting: Assume that in the parents’ cohort $c$ , $P$ is normally distributed with mean $p_{c}=2.0$, with standard deviation $\sigma_{c}=0.2$ The number of points in a genotype is $k=4$; the points in a genotype set are shown here in a sorted order for easier reading. The standard deviation of the normal distribution of the reproduction noise is $h=0.15$
@@
    Let individual $e_{2}$ in the parent cohort $c$ have genotype encoding of $g_{2}~{}=\{1.95,2.15,2.25,2.35\}$. whose mean and phenotype is $p_{2}=2.175$.
@@ 4 Comparing the Two Reproduction Methods
In comparing MR and BR in a particular transition from a parent cohort $c$ to a child cohort $c^{\prime}$, we are initially interested in a basic metric $d$ defined as the distance between the phenotype properties of a random individual in $c^{\prime}$, denoted $p^{\prime}_{i}$, and the common property $p_{c}$ of the parents’ cohort:
$d_{i}=|p^{\prime}_{i}-p_{c}|~{}.$
For a given parent cohort $c$, let $d_{\text{MR}}$ and $d_{\text{BR}}$ be the expected values $E[d_{i}]$ under the MR and BR processes, respectively, when considering all possible $c\shortrightarrow c^{\prime}$ transitions that could emanate from $c$ in one reproduction step, and random choices of $e^{\prime}_{i}$ individuals within $c^{\prime}$.
For a given parent cohort $c$, let $\sigma^{2}_{\text{MR}}$ and $\sigma^{2}_{\text{BR}}$ be the expected values $E[\sigma^{2}_{c^{\prime}}]$, the variance of $c^{\prime}$ under the MR and BR processes, respectively, when considering all possible $c\shortrightarrow c^{\prime}$ transitions that could emanate from $c$ in one reproduction step under the respective reproduction process.
@@
Then, $d_{\text{BR}}<d_{\text{MR}}$.
@@
A. Compute $d_{\text{MR}}$.
Let $c^{\prime}_{\text{MR}}$ be a random child cohort of $c$, selected from all possible MR transitions with $c$ as a parent.
Let $e^{\prime}_{i_{\text{MR}}}$ be a random individual in $c^{\prime}_{\text{MR}}$.
Let $g^{\prime}_{i_{\text{MR}}}$ be the genotype encoding of $e^{\prime}_{i_{\text{MR}}}$ and let $p^{\prime}_{i_{\text{MR}}}$ be its phenotype.
Let $e_{i}$ be the element of $c$ that served as the single parent of $e^{\prime}_{i_{\text{MR}}}$ in this MR transition.
@@
The genotype encodings $g^{\prime}_{i_{\text{MR}}}$ of all possible direct MR children of $e_{i}$, can be considered as samples of size $k$ from a normal distribution with mean $p_{i}$ and variance $h^{2}$, where $h$ is the noise parameter defined above.
@@
According to the central limit theorem [Blitzstein and Hwang, 2015, p.435], the means of these samples, namely $p^{\prime}_{i_{\text{MR}}}$, are distributed normally with mean $p_{i}$ and standard deviation $\frac{h}{\sqrt{k}}$ (and variance $\frac{h^{2}}{k}$).
@@
We are facing now a compound, nested distribution, that involves (i) the random selection of an $e_{i}$, which is close enough for our purposes to drawing a random value $p_{i}$ from the original normal distribution of $c$ with mean $p_{c}$ and variance $\sigma^{2}_{c}$, and (ii) based on the selected $p_{i}$ and the noise parameter, randomly selecting the sample $g^{\prime}$ of size $k$ and computing its mean $p^{\prime}_{i_{\text{MR}}}$.
@@
We can now examine the distribution of the random variable $p^{\prime}_{i_{\text{MR}}}$ as drawn from the original distribution of $c$. According to the law of total variance (a.k.a. Eve’s law) [Blitzstein and Hwang, 2015, Sec.9.5], this is a normal distribution with mean $p_{c}$ and variance that is the sum of the variances, hence
$\sigma^{2}_{c^{\prime}_{\text{MR}}}=\sigma^{2}_{c}+\frac{h^{2}}{k}$
@@
In general, the average absolute deviation of a random sample from the mean of a normal distribution with standard deviation $\sigma$ is $\sigma\cdot\sqrt{\frac{2}{\pi}}~{}~{}$  [Geary, 1935]. Therefore, the expected value of the absolute distance of $p^{\prime}_{i_{\text{MR}}}$ from $p_{c}$ is
@@
$d_{\text{MR}}=E[|p^{\prime}_{i_{\text{MR}}}-p_{c}|]=\left(\sqrt{\sigma^{2}_{c}+\frac{h^{2}}{k}}\right)\cdot\sqrt{\frac{2}{\pi}}~{}~{}.$
B. Compute $d_{\text{BR}}$.
@@
We now examine the expected effects of all possible BR transitions with $c$ as a parent cohort.
@@
Let $c^{\prime}_{\text{BR}}$ be a random child cohort of $c$ within all possible BR transitions.
Let $e^{\prime}_{j_{\text{BR}}}$ be a random individual in $c^{\prime}_{\text{BR}}$.
Let $g^{\prime}_{j_{\text{BR}}}$ be the genotype encoding of $e^{\prime}_{j_{\text{BR}}}$ and and let $p^{\prime}_{j_{\text{BR}}}$ be its phenotype.
@@
Let $e_{u}$ and $e_{v}$ be the two elements of $c$ that served as the parents of $e^{\prime}_{j_{\text{BR}}}$ in this BR transition, and let $p_{u}$ and $p_{v}$ be the means of their genotype encodings, respectively.
@@
The genotype encoding $g^{\prime}_{j_{\text{BR}}}$ was created as a union of two random samples $g^{\prime}_{u}$ and $g^{\prime}_{v}$, each of size $\frac{k}{2}$, around the means $p_{u}$ and $p_{v}$, respectively. Let $p^{\prime}_{u}$ and $p^{\prime}_{v}$ be the means of $g^{\prime}_{u}$ and $g^{\prime}_{v}$, respectively.
@@
The phenotype $p^{\prime}_{j_{\text{BR}}}$ is the mean of the union of $g^{\prime}_{u}$ and $g^{\prime}_{v}$, and is thus the average of the means of these two equal size samples:
$p^{\prime}_{j_{\text{BR}}}=\frac{p^{\prime}_{u}+p^{\prime}_{v}}{2}$
@@
Let us now analyze the random selection of $p^{\prime}_{j_{\text{BR}}}$ compounding the randomness inherent in the way $g^{\prime}_{u}$ and $g^{\prime}_{v}$ were sampled, and the nested randomness of $p_{u}$ and $p_{v}$ as selected from the original distribution of $c$.
@@
Compounding each of these two distributions separately with the underlying distribution of $c$ we get that $p^{\prime}_{u}$ and $p^{\prime}_{v}$ are distributed normally with mean $p_{c}$ and variance $\sigma^{2}_{c}+\frac{h^{2}}{\frac{k}{2}}$
Clearly the mean of the random variable $p^{\prime}_{j_{\text{BR}}}$ is $p_{c}$ since it is half of the sum $E[p^{\prime}_{u}]$ and $E[p^{\prime}_{v}]$ (each of which is equal to $p_{c}$), and its variance is a quarter of the variance of their sum, as it is the square of the scaling by half of the variable and the mean [Blitzstein and Hwang, 2015, p.159]:
$\sigma^{2}_{\text{BR}}=\frac{1}{4}\cdot 2(\sigma^{2}_{c}+\frac{h^{2}}{\frac{k}{2}})=\frac{\sigma^{2}_{c}}{2}+\frac{h^{2}}{k}.$
As before,
$d_{\text{BR}}=E[|p^{\prime}_{j_{\text{BR}}}-p_{c}|]=\left(\sqrt{\frac{\sigma^{2}_{c}}{2}+\frac{h^{2}}{k}}\right)\cdot\sqrt{\frac{2}{\pi}}~{}~{}.$
C. Comparing $d_{\text{BR}}$ and $d_{\text{MR}}$.
Since the first term in the expression for $d_{\text{BR}}$ is $\frac{\sigma^{2}_{c}}{2}$ and for $d_{\text{MR}}$ it is $\sigma^{2}_{c}$, we get that
$d_{\text{BR}}<d_{\text{MR}}~{}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\square$
Note that an intermediate result in the above calculations is that $\sigma^{2}_{\text{BR}}<\sigma^{2}_{\text{MR}}$. This aligns well with the established fact, as was discussed earlier, that in nature, sexual reproduction yields a lower variance within the new generation than does mono-parental reproduction. In the present context, such smaller values of $\sigma_{c^{\prime}}$ help individuals in $c^{\prime}$ step in and replace each other when needed, independently of the magnitude of their differences from their parents and from earlier generations.

### v5

@@ 3.2 Reproduction methods
Below we describe one reproduction step of an entire generation, termed a generational transition, in each of the methods.
Mono-parental Reproduction (MR)
@@
    The input is a set $c$ of the current population; create an empty set $c^{\prime}$.
@@
      $p_{i}=\frac{1}{k}\sum_{\ell=1}^{k}x_{i_{\ell}}\leavevmode\nobreak\ \leavevmode\nobreak\ .$
@@
    After completing $n$ iterations: discard the parent set $c$; rename the child set $c^{\prime}$ to $c$; discard $c^{\prime}$.
Bi-parental Reproduction (BR)
@@
    As in MR, the input is a set $c$ of the current population; create an empty set $c^{\prime}$.
@@
      Pick a random pair of two separate individuals, $e_{u},e_{v}\in c,\leavevmode\nobreak\ u\neq v,1\leq u,v\leq n$ out of the $\frac{n(n-1)}{2}$ possible pairs in $c$. The pair $e_{u},e_{v}$ will be the joint parents of the child created in this iteration; in separate iterations these two individuals may participate in other pairs or again in the very same pair.
@@
    As in MR, after completing $n$ iterations: discard the parent set $c$; rename the child set $c^{\prime}$ to $c$; discard $c^{\prime}$.
As stated earlier, neither MR nor BR employ considerations of Natural Selection and classical fitness metrics like an individual’s probability of reproduction and its expected number of offspring over its life time, and BR does not make use of mating types or mate selection.
@@ 3.3 Model realization example
Below we illustrate the above formal model with an example realization in an imaginary species of intelligent and capable animals. One of the traits that sustains this species is their ability to build every year a mud hill of “exactly” a certain height, as traps for some creature that they feed upon. The way they remember the right height is that each mature individual carries with them a bunch of sticks that they got from their parents. The individual builds each mud hill so that its height is close to the mean length of its set of sticks. Old mud hills from the previous year are not available for comparison, as they have been washed away by the rains. When young individuals mature, they leave their parents to look for new territory. Before they leave, the parents (one or two of them, as the case may be) equip the young with a new bunch of sticks, cut as best as they can to the mean length of the parents’ bunch. In the case of two parents, each one gives the child half the number of sticks based on that parent’s set. The parents cannot give their own sticks to the offspring, as they have to keep their own sticks for several more years. The individuals are solitary most of the time, so each one needs their own set of sticks. Individuals of the species use sets of sticks for this purpose, rather than just one, as the sticks may break or get lost. In addition, the parent’s skill and the available material limit the precision of each cut stick; having several sticks, some too short and some too long, is a way for conveying the height of the mud hill that is desired by the individual preparing the new set.
@@ 3.4 A numerical example
    Parameter Setting: Assume that in the parents’ cohort $c$ , $P$ is normally distributed in the parent generation with mean $p_{c}=2.0$, and with standard deviation $\sigma_{c}=0.2$ The number of points in a genotype is $k=4$; the points in a genotype set are shown here in a sorted order for easier reading. The standard deviation of the normal distribution of the reproduction noise is $h=0.15$;
@@
    Let individual $e_{2}$ in the parent cohort $c$ have genotype encoding of $g_{2}\leavevmode\nobreak\ =\{1.95,2.15,2.25,2.35\}$, whose mean and phenotype is $p_{2}=2.175$.
@@ 4 Comparing the Two Reproduction Methods
In comparing MR and BR in a particular generational transition from a parent cohort $c$ to a child cohort $c^{\prime}$, we are initially interested in a basic metric $d$ defined as the distance between the phenotype trait of a random individual in $c^{\prime}$, denoted $p^{\prime}_{i}$, and the common trait $p_{c}$ of the parents’ cohort:
$d_{i}=|p^{\prime}_{i}-p_{c}|\leavevmode\nobreak\ .$
For a given parent cohort $c$, let $d_{\text{\tiny MR}}$ and $d_{\text{\tiny BR}}$ be the expected values $E[d_{i}]$ under the MR and BR processes, respectively, when considering all possible $c\shortrightarrow c^{\prime}$ generational transitions that could emanate from $c$ in one reproduction step, and random choices of $e^{\prime}_{i}$ individuals within $c^{\prime}$.
For a given parent cohort $c$, let $\sigma^{2}_{\text{\tiny MR}}$ and $\sigma^{2}_{\text{\tiny BR}}$ be the expected values $E[\sigma^{2}_{c^{\prime}}]$, the variance of $P$ within $c^{\prime}$ under the MR and BR processes, respectively, when considering all possible $c\shortrightarrow c^{\prime}$ generational transitions that could emanate from $c$ in one reproduction step under the respective reproduction process.
@@
Then, $d_{\text{\tiny BR}}<d_{\text{\tiny MR}}$.
@@
A. Compute $d_{\text{\tiny MR}}$.
Let $c^{\prime}_{\text{\tiny MR}}$ be a random child cohort of $c$, selected from all possible MR generational transitions with $c$ as a parent.
Let $e^{\prime}_{i_{\text{\tiny MR}}}$ be a random individual in $c^{\prime}_{\text{\tiny MR}}$.
Let $g^{\prime}_{i_{\text{\tiny MR}}}$ be the genotype encoding of $e^{\prime}_{i_{\text{\tiny MR}}}$, and let $p^{\prime}_{i_{\text{\tiny MR}}}$ be its phenotype.
Let $e_{i}$ be the element of $c$ that served as the single parent of $e^{\prime}_{i_{\text{\tiny MR}}}$ in this MR generational transition.
@@
The genotype encodings $g^{\prime}_{i_{\text{\tiny MR}}}$ of all possible direct MR children of $e_{i}$, can be considered as samples of size $k$ from a normal distribution with mean $p_{i}$ and variance $h^{2}$, where $h$ is the noise parameter defined above.
@@
According to the central limit theorem [27](p.435), the means of these samples, namely $p^{\prime}_{i_{\text{\tiny MR}}}$, are distributed normally with mean $p_{i}$ and standard deviation $\frac{h}{\sqrt{k}}$ (and variance $\frac{h^{2}}{k}$).
@@
We are facing now a compound, nested distribution, that involves (i) the random selection of an $e_{i}$, which is close enough for our purposes to drawing a random value $p_{i}$ from the original normal distribution of $c$ with mean $p_{c}$ and variance $\sigma^{2}_{c}$, and (ii) based on the selected $p_{i}$ and the noise parameter, randomly selecting the sample $g^{\prime}$ of size $k$ and computing its mean $p^{\prime}_{i_{\text{\tiny MR}}}$.
@@
We can now examine the distribution of the random variable $p^{\prime}_{i_{\text{\tiny MR}}}$ as drawn from the original distribution of $c$. According to the law of total variance (a.k.a. Eve’s law [27](p.401)), this is a normal distribution with mean $p_{c}$ and variance that is the sum of the variances, hence
$\sigma^{2}_{c^{\prime}_{\text{\tiny MR}}}=\sigma^{2}_{c}+\frac{h^{2}}{k}$
@@
In general, the average absolute deviation of a random sample from the mean of a normal distribution with standard deviation $\sigma$ is $\sigma\cdot\sqrt{\frac{2}{\pi}}\leavevmode\nobreak\ \leavevmode\nobreak$ [28]. Therefore, the expected value of the absolute distance of $p^{\prime}_{i_{\text{\tiny MR}}}$ from $p_{c}$ is
@@
$d_{\text{\tiny MR}}=E[\leavevmode\nobreak\ |p^{\prime}_{i_{\text{\tiny MR}}}-p_{c}|\leavevmode\nobreak\ ]=\left(\sqrt{\sigma^{2}_{c}+\frac{h^{2}}{k}}\right)\cdot\sqrt{\frac{2}{\pi}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ .$
B. Compute $d_{\text{\tiny BR}}$.
@@
We now examine the expected effects of all possible BR generational transitions with $c$ as a parent cohort.
@@
Let $c^{\prime}_{\text{\tiny BR}}$ be a random child cohort of $c$ within all possible BR generational transitions.
Let $e^{\prime}_{j_{\text{\tiny BR}}}$ be a random individual in $c^{\prime}_{\text{\tiny BR}}$.
Let $g^{\prime}_{j_{\text{\tiny BR}}}$ be the genotype encoding of $e^{\prime}_{j_{\text{\tiny BR}}}$ and and let $p^{\prime}_{j_{\text{\tiny BR}}}$ be its phenotype.
@@
Let $e_{u}$ and $e_{v}$ be the two elements of $c$ that served as the parents of $e^{\prime}_{j_{\text{\tiny BR}}}$ in this BR generational transition, and let $p_{u}$ and $p_{v}$ be the means of their genotype encodings, respectively.
@@
The genotype encoding $g^{\prime}_{j_{\text{\tiny BR}}}$ was created as a union of two random samples $g^{\prime}_{u}$ and $g^{\prime}_{v}$, each of size $\frac{k}{2}$, around the means $p_{u}$ and $p_{v}$, respectively. Let $p^{\prime}_{u}$ and $p^{\prime}_{v}$ be the means of $g^{\prime}_{u}$ and $g^{\prime}_{v}$, respectively.
@@
The phenotype $p^{\prime}_{j_{\text{\tiny BR}}}$ is the mean of the union of $g^{\prime}_{u}$ and $g^{\prime}_{v}$, and is thus the average of the means of these two equal size samples:
$p^{\prime}_{j_{\text{\tiny BR}}}=\frac{p^{\prime}_{u}+p^{\prime}_{v}}{2}$
@@
Let us now analyze the random selection of $p^{\prime}_{j_{\text{\tiny BR}}}$ compounding the randomness inherent in the way $g^{\prime}_{u}$ and $g^{\prime}_{v}$ were sampled, and the nested randomness of $p_{u}$ and $p_{v}$ as selected from the original distribution of $c$.
@@
Compounding each of these two distributions separately with the underlying distribution of $c$ we get that $p^{\prime}_{u}$ and $p^{\prime}_{v}$ are distributed normally with mean $p_{c}$ and variance that is the sum $\sigma^{2}_{c}+\frac{h^{2}}{\frac{k}{2}}$
Clearly the mean of the random variable $p^{\prime}_{j_{\text{\tiny BR}}}$ is $p_{c}$ since it is half of the sum $E[p^{\prime}_{u}]$ and $E[p^{\prime}_{v}]$ (each of which is equal to $p_{c}$), and its variance is a quarter of the variance of their sum, as it is the square of the scaling by half of the variable and the mean [27](p.159):
$\sigma^{2}_{\text{\tiny BR}}=\frac{1}{4}\cdot 2(\sigma^{2}_{c}+\frac{h^{2}}{\frac{k}{2}})=\frac{\sigma^{2}_{c}}{2}+\frac{h^{2}}{k}\leavevmode\nobreak\ .$
As before, the average absolute deviation is given by
$d_{\text{\tiny BR}}=E[|p^{\prime}_{j_{\text{\tiny BR}}}-p_{c}|]=\left(\sqrt{\frac{\sigma^{2}_{c}}{2}+\frac{h^{2}}{k}}\right)\cdot\sqrt{\frac{2}{\pi}}\leavevmode\nobreak\ \leavevmode\nobreak\ .$
C. Comparing $d_{\text{\tiny BR}}$ and $d_{\text{\tiny MR}}$.
The expressions for $d_{\text{\tiny MR}}$ and $d_{\text{\tiny BR}}$ are the same except for the first term in $d_{\text{\tiny BR}}$ being $\frac{\sigma^{2}_{c}}{2}$ where in $d_{\text{\tiny MR}}$ it is $\sigma^{2}_{c}$. Hence,
$d_{\text{\tiny BR}}<d_{\text{\tiny MR}}\leavevmode\nobreak\ .\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \square$
Note that an intermediate result in the above calculations is that $\sigma^{2}_{\text{\tiny BR}}<\sigma^{2}_{\text{\tiny MR}}$. This aligns well with the established fact, as was discussed earlier, that in nature, sexual reproduction yields a lower variance within the new generation than does mono-parental reproduction. In the present context, such smaller values of $\sigma_{c^{\prime}}$ help individuals in $c^{\prime}$ step in and replace each other when needed, independently of the magnitude of their differences from their parents and from earlier generations.